OpenStudy (anonymous):

e^x - e^-x/e^x+e^-x = 1/2

OpenStudy (henryblah):

\[(e^x-e ^{-x})/(e^x+e ^{-x})=0.5\]

OpenStudy (anonymous):

Left hand side = tanh x (hyperbolic tangent)

OpenStudy (anonymous):

@Henryblah how do I solve the rest?

OpenStudy (anonymous):

You can rewrite tanh x as e^(2x)-1 / e^(2x) +1 Multiply out, simplify then take logs.

OpenStudy (anonymous):

@estudier where did you get 2x from?

OpenStudy (anonymous):

e^x + e^-x ie e^x + 1/e^x Add them

OpenStudy (anonymous):

A little more assistance here please.

OpenStudy (lgbasallote):

\(\LARGE e^x + \frac{1}{e^x} =\frac{e^{2x} + 1}{e^x}\quad \text{make sense?}\)

OpenStudy (lgbasallote):

that's the denominator ^

OpenStudy (lgbasallote):

\(\Large e^x - \frac{1}{e^x} = \frac{e^{2x} - 1}{e^x} \leftarrow \text{numerator}\)

OpenStudy (lgbasallote):

@math102 are you understanding so far?

OpenStudy (anonymous):

So we are now subtracting 1/e^x because I see now you have -1/e^x

OpenStudy (lgbasallote):

yepp

OpenStudy (lgbasallote):

so if we rewrite \(\Large \frac{e^x - e^{-x}}{e^x + e^{-x}}\) you'll have \[\Huge \frac{\frac{e^{2x} - 1}{e^x}}{\frac{e^{2x} + 1}{e^x}}\] right?

OpenStudy (anonymous):

Okay.

OpenStudy (lgbasallote):

cancel the \(\large e^x\) on both doth denominators and you'll get to where @estudier left off :D

OpenStudy (anonymous):

Keep going...

OpenStudy (lgbasallote):

uhmm well now we have \(\LARGE \frac{e^{2x} - 1}{e^{2x} + 1} = \frac{1}{2}\) if we cross multiply... \(\LARGE 2e^{2x} - 2 = e^{2x} + 1\) do you know how to go from there?

OpenStudy (anonymous):

Okay so now I have e^2x -1/ e^2x + 1

OpenStudy (lgbasallote):

yup then that \(\uparrow\)

OpenStudy (anonymous):

\[e ^{2x}-1/ e ^{2x}+1\]

OpenStudy (anonymous):

Okay so now what do I do with the 2x's?

OpenStudy (lgbasallote):

nothing yet..continuing from what i said earlier..you will put the e^2x on one side and the constants on the opposite

OpenStudy (anonymous):

So 2e^2x=2-1?

OpenStudy (lgbasallote):

2e^2x - e^2x = 2 + 1

OpenStudy (anonymous):

So I have 2e^2x-e^2x=3

OpenStudy (lgbasallote):

and what is 2e^2x - e^2x?

OpenStudy (anonymous):

So I dont use the 3 on the right side of the equation?

OpenStudy (lgbasallote):

no..there's a 3..im asking you to simplify 2e^2x - e^2x

OpenStudy (anonymous):

Would it be 1e^2x?

OpenStudy (lgbasallote):

yup! but 1 is no longer needed to write so we have e^2x = 3 take the ln of both sides

OpenStudy (anonymous):

How do I take the natural log of e^2x?

OpenStudy (lgbasallote):

let's put it as ln e^2x for now

OpenStudy (anonymous):

So e ln 2?

OpenStudy (lgbasallote):

uhmm no...wait.. you still have equal 3 ln e^2x = ln 3 do a power rule on ln e^2x

OpenStudy (anonymous):

I tbink I get it now 2x= ln 3 x=ln 3/2 x=.5493?

OpenStudy (anonymous):

After e^2x=3 should I have written 2x ln e =3?

OpenStudy (lgbasallote):

uhmm lemme write it ln e^2x = ln 3 2x ln e = ln 3 2x (1) = ln 3 x = (ln 3)/2

OpenStudy (anonymous):

What happened to the 2x(1) in the third step you wrote?

OpenStudy (anonymous):

Nevermind I got it. Thank you so much. You've been more than helpful. Superb!

OpenStudy (lgbasallote):

no biggie :D